OpenStudy (anonymous):

please help me how to integrate x^2d(x^2)..? and x^3d(x^2)...

OpenStudy (shubhamsrg):

in first one,,let x^2=t and you are left with integral t dt.. in second one,,,let x^2 =t,,so x^3 = t^3/2 //so you need to integrate t^3/2 dt.. should be easier then..

OpenStudy (anonymous):

I have done earlier as you told but mine answer was not matched, then I post the problem, there might be answer -mistake in book.

OpenStudy (shubhamsrg):

(x^4)/2 and (x^5) * 2/5 you sure these dont match ?